# p-MoD: Building Mixture-of-Depths MLLMs via Progressive Ratio Decay

### v1

@@ 3.3 Progressive Ratio Decay
GQA RWQA AI2D MMMU AVG
LLaVA-NeXT 560 39.46 5191 100% 61.6 34.7 63.5 57.6 64.0 35.3 52.8
p-MoD-0.3 403 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-28.0\%)}$ 17.78 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-54.9\%)}$ 3267 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-37.1\%)}$ 42.3% 57.8 32.1 63.1 55.9 65.0 35.0 51.5
p-MoD-0.4 415 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-25.9\%)}$ 19.56 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-50.4\%)}$ 3471 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-33.1\%)}$ 47.5% 59.6 33.0 63.8 57.3 65.0 35.1 52.3
p-MoD-0.5 435 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-22.3\%)}$ 21.94 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-44.4\%)}$ 3680 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-29.1\%)}$ 53.8% 61.8 34.1 63.3 57.6 65.1 36.0 53.0
In addition, we find it challenging for MoD layers to learn to predict meaningful weights when the token retention ratio is set to an extreme value. To address this issue we constrain the range of the token retention ratio $R_{l}$ within a predefined maximum and minimum threshold. If $R_{l}$ exceeds the maximum threshold, we set the token retention ratio to $1$ so that the MoD module is not applied to the $l$-th layer. If $R_{l}$ falls below the minimum threshold, the token retention ratio for the $l$-th layer is set to the minimum threshold:
$R_{l}^{\prime}=\begin{cases}1,&\text{if}\quad R_{l}\geq max\\ R_{l},&\text{if}\quad min<R_{l}<max\\ min,&\text{if}\quad R_{l}\leq min\end{cases}$ (6)
4 Experiment
4.1 Setups
Models. To evaluate the effectiveness of p-MoD, we select two representative open-source MLLMs: LLaVA-1.5 [23] and LLaVA-NeXT [24], as the baseline models in our experiments. We name our corresponding p-MoD models p-MoD-LLaVA-1.5 and p-MoD-LLaVA-NeXT. LLaVA-1.5 resizes input images to the fixed resolution which aligns with the training setup of its CLIP [34] vision encoder, encoding an image into 576 tokens. LLaVA-NeXT divides high-resolution images into multiple slices which are independently processed by the vision encoder. This strategy enhances the visual perception capabilities, but also leads to a significantly larger number of vision tokens (up to 2880) and much higher computation costs.
Training Data. We follow the training setups of our baseline methods. Both LLaVA-1.5 and LLaVA-NeXT utilize LCS-558K [25] as pre-training data. During the instruction tuning stage, LLaVA-1.5 utilizes LLaVA-mix-665k[23]. The original LLaVA-NeXT model utilizes 760k data for fine-tuning, however, the LLaVA-NeXT training data released by LLaVA team contains 779k data. For fair comparison, we train LLaVA-NeXT model on the released 779k data with the official LLaVA-NeXT code base.
Benchmarks. We conduct comprehensive experiments across 14 benchmarks: GQA [15], OK-VQA [29], AI2D [17] and ScienceQA-IMG [28] are traditional visual question answering benchmarks; DocVQA [31], TextVQA [37], ChartQA [30] and InfographicVQA [32] focus on fine-grained visual question answering; POPE [22] evaluates hallucination in MLLMs; SEED-Bench (Image) [19], RealWorldQA, MME[11], MMBench [26] and MMMU  [42] are comprehensive benchmarks tailored for MLLMs. To ensure our results can be conveniently reproduced, we evaluate our model on all these benchmarks with the lmms-eval [43] evaluation framework developed and maintained by LLaVA team.
Evaluation Metrics. We report the performance on all benchmarks, together with a variety of metrics that reflect training and inference efficiency. We report TFLOPs which measures the inference computation complexity, and the KV cache storage which is the main memory bottleneck during inference. Besides, we also report the GPU hours during model training, and the inference latency.
Implementation Details. We train all our models on 8 NVIDIA RTX A6000 GPUs. The inference efficiency metrics reported in Section 4.3 are measured on a single A6000 GPU. By default, we set the gating factor $\alpha$ in TanhNorm to 0.2, and the shift factor $\beta$ in PRD to $0.5$.
4.2 Main Results
We integrate p-MoD with LLaVA-1.5 and LLaVA-NeXT baseline models and conduct comprehensive evaluation across 14 benchmarks, as shown in table 1 and table 2. Both p-MoD-LLaVA-v1.5 and p-MoD-LLaVA-NeXT achieve comparable or even better performance across all benchmarks compared to their baseline models, with significant savings in inference TFLOPs and inference KV cache savings.
Table 1 consists of general VQA and MLLM benchmarks, which measures comprehensive abilities of MLLMs across various aspects. Our model p-MoD-LLaVA-NeXT model achieves comparable performance while significantly reducing inference TFLOPs and KV cache storage. Our p-MoD-LLaVA-v1.5 model even outperforms the LLaVA-1.5 baseline model, with only 58.7% TFLOPs and 53.8% KV cache storage.
4.3 Efficiency-Performance Trade-off
To showcase the versatility of our approach, we conducted experiments with $\beta=0.3,0.4,0.5$, respectively. The results are presented in Table 3. Our default p-MoD model with $\beta=0.5$ reduced training GPU hours by 22.3%, inference TFLOPs by 44.4%, inference latency by 29.1%, and KV cache storage by 46.2%. Using a lower $\beta$ can further improve efficiency, with an acceptable performance drop.
In conclusion, Table 3 demonstrates a trend that as the computation cost is reduced, p-MoD’s performance drops linearly with a shallow slope, offering a superior performance-efficiency trade-off.
Norm
Type
STR-
ing
Doc
VQA
@@ 4.3 Efficiency-Performance Trade-off
SEED RWQA AI2D MMB AVG
[…]
Table 4: Ablation on tanh-gated weight normalization and symmetric token reweighting. Our default model is marked in gray. The rows with no performance reported indicates that the corresponding experiment faces overflow issue during training or inference.
Ratio
Schedule
Decay Progressive
Doc
@@
SEED RWQA AI2D MMB AVG
[…]
4.4 Ablation Study
Tanh-gated Weight Normalization. To validate the effectiveness of TanhNorm, we explore the effectiveness of different weight normalization methods in Table 4. First, we verify that using no weight normalization module will result in overflow during training, as shown in the first row. Then, we validate the effectiveness of the two important properties of TanhNorm stated in Section 3.2.1: (1) the normalized weight distribution is zero-centered, (2) the range(variance) of the weight distribution can be easily controlled by adjusting the gating factor $\alpha$.
Figure 5: Visualization of tokens selected by different p-MoD layers. The selected tokens are colored in red. We observe that vision tokens corresponding to regions with rich semantic information are selected across different p-MoD layers, even when the number of selected tokens becomes minimal in deeper layers.
Symmetric Token Reweighting. By comparing the last two rows in Table 4, we can confirm that employing the symmetric token reweighting(STRing) strategy leads to a 1% improvement on both DocVQA and RWQA, and the average performance is enhanced by 0.3%. It is verified that STRing successfully improves performance by more effectively leveraging language supervision signals to train a superior weight predictor with limited training data.
To further validate this conclusion and the effectiveness of PRD, we experiment with four other schedule functions which control the token retention ratio for each decoder layer. As illustrated in Figure 4, the functions include: (1) a constant function that uses the same ratio for all layers, (2) an interleaved function which interleaves a vanilla transformer layer and an MoD layer with low ratio, which is adopted in the original MoD paper [35], (3) a stepped decay function, (4) a linear decay function. To ensure a fair comparison, we fix the average retention ratios across all layers at approximately 54%. Based on the results demonstrated in Table 5, we can draw the following conclusions:
  • The continuous decay functions (i.e. cosine and linear) notably outperforms the discontinuous stepped decay function, and the cosine function achieves the best results. This proves the effectiveness of our shifted cosine schedule.
4.5 Visualization
Figure 5 depicts the tokens selected by p-MoD-LLaVA-NeXT at different layers. Guided by the progressive ratio decay strategy, the number of tokens processed by the MoD layers gradually decreases layer by layer. It can be observed that the selected vision tokens progressively concentrate on key regions in the image with rich semantic information. Specifically, in the figure above, visual tokens corresponding to text and drawings in different positions are nearly all selected at the 26th layer with a low ratio. In the figure below, selected tokens gradually converge towards measurement markings on a ruler and the other object. These results indicate that our model effectively compresses visual information by keeping the informative tokens, enabling it to maintain performance while significantly reducing training and inference costs.
5 Conclusion and Future Work
In this paper, we have explored building efficient MLLMs with Mixture-of-Depths mechanism. Our proposed model, termed p-MoD, features three key designs. First, tanh-gated weight normalization mechanism is applied to the predicted token weights to ensure proper training initialization and stability during training and inference. Second, a symmetric token reweighting strategy is utilized to fully leverage the language supervision signals to train a superior weight prediction module with limited training data. Finally, our model employs a progressive ratio decay strategy to gradually reduce the token retention ratio layer by layer. Our p-MoD model achieves comparable or even superior results to the baseline models on a diverse set of 14 benchmarks, with substantial improvements on training and inference efficiency. We hope p-MoD can serve as a strong and efficient baseline for future research on developing efficient MLLMs.
One limitation of our work is that p-MoD is only experimented on LLaVA-1.5 and LLaVA-NeXT models, which focus on single-image understanding tasks. We believe that p-MoD has the potential to achieve more remarkable results when applied on tasks that handle a larger number of vision tokens, such as high-resolution, multi-image, and long video understanding. We leave the exploration of p-MoD on other vision tasks to future research.
Appendix
This supplementary material includes the following sections:
  • In Section A, we compare p-MoD against pruning methods that drop vision tokens using attention scores received from text tokens.
  • In Section B, we provide visualization results of p-MoD token selection decisions across different layers.
  • In Section C, we provide more implementation details.
Appendix A Comparison with Attention-Score-Based Pruning Methods
In this section, we compare p-MoD against pruning methods that drop vision tokens in intermediate transformer layers according to the attention scores they received from text tokens [4, 45, 14, 12]. As mentioned in Section 2, these methods are the closest to our work. First, we provide a detailed analysis on their limitations. Then, we conduct experiments to support our analysis.
A.1 Limitations of Attention-Score-Based Pruning Methods
Methods that use attention scores to drop vision tokens have the following limitations:
  • Incompatible with Efficient Attention Implementations. These methods require the attention module to output the attention score matrix to perform token dropping. However, this is incompatible with efficient attention implementations like FlashAttention [9, 8] which uses tiling to avoid materialization of the large attention score matrix on GPU HBM. Consequently, these methods can only utilize the inefficient vanilla attention implementation and fail to achieve the theoretically expected memory and time efficiency, despite the reduction in vision tokens. In contrast, our method imposes no restrictions on the implementation of attention mechanism and is therefore compatible with any attention implementation.
  • Dropping Tokens is Inferior to Selecting Tokens. Attention-score-based pruning methods drop vision tokens in certain transformer layers, which means the dropped tokens are not available in later layers. Intuitively, this strategy hurts performance in scenarios where some tokens with low attention scores in earlier layers are dropped, but they are crucial (i.e. should have received high attention scores) in later layers. Alternatively, our p-MoD model selects important tokens to process from the complete token set in every layer, mitigating the potential harm caused by token dropping.
A.2 Experiments
We conduct experiments on p-MoD and FastV [4], the most representative attention-score-based pruning method. FastV drops 50% vision tokens with the lowest attention scores from the generated text tokens, after the second transformer decoder layer. This setting ensures approximately the same TFLOPs as p-MoD. We choose LLaVA-NeXT as baseline model, and enable KV cache during generation for fair comparison. With KV cache enabled, FastV drops vision tokens according to the attention scores from the first generated text token.
The results are illustrated in Table 6. Since FastV is incompatible with FlashAttention, FastV-LLaVA-NeXT model can only utilize the vanilla attention implementation, which even consumes more GPU memory than the baseline model. Furthermore, FastV’s token dropping strategy results in significant performance degradation. In contrast, our p-MoD-LLaVA-NeXT model matches the performance of the baseline model, with the lowest TFLOPs, latency and memory usage during inference. This highlights the benefits of p-MoD’s token selection mechanism combined with its compatibility with efficient attention implementations.
@@ A.2 Experiments
Appendix B Visualization of p-MoD Token Selection Decisions
@@
Figure 6 visualizes the token selection decisions of p-MoD across different layers. The horizontal axis denotes the token indexes, and the vertical axis denotes the layer indexes. It can be observed that every layer select different tokens to process, and every token is selected by different p-MoD layers. This demonstrates that every p-MoD layer independently selects important and informative tokens, without degrading into selecting a same set of tokens across different layers, which is identical to dropping tokens instead of selecting.
Appendix C More Implementation Details
C.1 Detailed Training Configurations

### v2

@@ 4.4 Comparison with Related Works
Text
VQA
G
QA
SE
ED
$M$
B
AVG
LLaVA-NeXT-7B 560 39.46 519.1 100% 70.1 61.6 62.7 63.5 68.9 67.5 65.7
p-MoD-0.3 403 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-28.0\%)$} 17.78 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-54.9\%)$} 326.7 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-37.1\%)$} 42.3% 65.5 57.8 58.1 63.1 68.6 67.0 63.4
p-MoD-0.4 415 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-25.9\%)$} 19.56 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-50.4\%)$} 347.1 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-33.1\%)$} 47.5% 66.5 59.6 58.4 63.8 68.5 66.6 63.9
p-MoD-0.5 435 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-22.3\%)$} 21.94 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-44.4\%)$} 368.0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-29.1\%)$} 53.7% 70.0 61.8 60.5 63.3 69.0 67.3 65.3
Figure 5: Visualization of tokens selected by different p-MoD layers. The selected tokens are colored in red. Please zoom in for a clearer view. As the number of selected tokens gradually decreases in deeper layers (due to PRD strategy), the model gradually concentrates on vision tokens corresponding to regions with rich semantic information.
4.5 Efficiency-Performance Trade-off
To showcase the versatility of our approach, we conducted experiments with $\beta=0.3,0.4,0.5$, respectively. The results are presented in Table 7. Our default p-MoD model with $\beta=0.5$ reduced training GPU hours by 22.3%, inference TFLOPs by 44.4%, inference latency by 29.1%, and KV cache storage by 46.3%. Using a lower $\beta$ can further improve efficiency, with an acceptable performance drop.
4.6 Visualization: Which tokens are selected?
In Figure 5, we visualize the tokens selected by p-MoD-LLaVA-NeXT at different layers. Guided by the PRD strategy, the number of tokens processed by the MoD layers gradually decreases layer by layer. It can be observed that the selected vision tokens progressively concentrate on key regions in the image with rich semantic information. For the upper row of images in Figure 5, the selected tokens gradually converge to those corresponding to text, drawings, and the person on the right. Tokens corresponding to the white background are skipped in deep layers. In the lower row of images in Figure 5, the selected tokens gradually converge towards the artifact and the measurement markings on the ruler. These results indicate that our model effectively compresses visual information by selecting the informative tokens, enabling it to maintain performance while significantly reducing training and inference costs.
5 Conclusion and Future Work
In this paper, we explore building efficient MLLMs by adapting the Mixture-of-Depths mechanism. Our proposed model, termed p-MoD, features three key designs: tanh-gated weight normalization, symmetric token reweighting module, and the progressive ratio decay strategy. Our model achieves comparable or even superior results to the baseline models on a diverse set of 15 benchmarks, with substantial improvements on training and inference efficiency. We hope p-MoD can serve as a strong and efficient baseline for future research on developing efficient MLLMs.
Acknowledgements. This work is supported by the National Key R&\&D Program of China (No. 2022ZD0160900), Jiangsu Frontier Technology Research and Development Program (No. BF2024076), the Collaborative Innovation Center of Novel Software Technology and Industrialization, and Nanjing University-China Mobile Communications Group Co., Ltd. Joint Institute.
Supplementary Material
This supplementary material includes the following sections:
  • In Section A, we provide more experiment results.
  • In Section B, we show visualization results of p-MoD token selection decisions across different layers.
  • In Section C, we make some further discussions about our work.
  • In Section D, we provide more implementation details.
Appendix A More Experiments
A.1 Experiments on Larger Model Size
In Table 8, we compare our method against other token compression methods on LLaVA-NeXT-13B baseline model. The results demonstrate that our method significantly outperforms other methods on larger model size, validating the scalability of our approach.
A.2 Experiments on Visual Grounding
In Table 9, we compare p-MoD with other token compression methods on visual grounding benchmarks RefCOCO, RefCOCO+ and RefCOCOg [20, 52]. Our method significantly outperforms other methods, but it still exhibits some performance drop compared to the baseline model. This suggests that current token pruning methods still exhibit limitations on visual grounding, which might be one of the major challenges that future works in this field should aim to address.
A.3 Compairson with More Relate Works
In addition to comparing p-MoD against three strong token compression methods in Table 5, we provide comparison with more methods in Table 10. Under the same token compression ratio, p-MoD demonstrates the best overall performance on our comprehensive evaluation suite covering 15 benchmarks.
A.4 Comparison with $\gamma$-MoD
Concurrent to our work, $\gamma$-MoD [33] also propose to integrate Mixture-of-Depths mechanism into MLLMs. In this section, we first analyze the difference between our approach and theirs. Then we conduct experiments to show that our p-MoD approach outperforms $\gamma$-MoD.
The core design of $\gamma$-MoD is computing attention map (ARank) on some samples to identify which layers MoD can be applied to. In contrast, p-MoD can be effectively applied to every layer thanks to our TanhNorm and STRing modules. Furthermore, we propose PRD strategy to progressively reduce the token retention ratio layer by layer, which significantly boosts performance and efficiency.
In Table 11, we compare our method with $\gamma$-MoD on LLaVA-v1.5-7B baseline. $\gamma$-MoD utilizes higher token retention ratio than p-MoD (60+% vs 53%), but p-MoD still outperforms it by a large margin. Note that we are unable to compare both methods on LLaVA-NeXT, as $\gamma$-MoD’s code implementation does not support LLaVA-NeXT.
Method DocVQA ChartQA SEED GQA
AVG over
15 Tasks*
 LLaVA-NeXT-13B 73.5 67.2 71.4 65.2 66.5
   + MQT 58.6 54.6 68.8 64.2 63.0
   + FastV 70.2 64.7 70.8 64.7 65.7
   + LLaVolta 70.0 61.7 70.5 64.7 65.0
   + p-MoD 72.3 66.2 71.6 65.0 66.0
Table 8: Experiments on LLaVA-NeXT-13B. Our method significantly outperforms other methods on larger baseline model, validating the scalability of our approach. *Average is computed on all 15 benchmarks used in Table 1 and 3.
Method RefCOCO Val RefCOCO+ Val RefCOCOg Val
LLaVA-NeXT-7B 82.67 73.73 79.41
+MQT 68.77 59.07 63.52
+LLaVolta 79.96 70.89 76.79
+FastV 73.83 64.90 69.78
+p-MoD 80.23 70.94 76.96
Table 9: Experiments on Visual Grounding. Our method significantly outperforms other methods on visual grounding benchmarks RefCOCO, RefCOCO+ and RefCOCOg. We report ACC@0.5 metric on the validation sets of these benchmarks.
Model DocVQA ChartQA SEED GQA
AVG over
15 tasks*
LLaVA-NeXT-7B 70.1 61.6 68.9 63.5 63.5
+ SparseVLM [57] 67.2 52.8 68.1 62.6 62.3
+ VisionZip [49] 65.5 50.3 67.4 61.1 61.2
+ PyramidDrop [48] 66.5 53.3 67.5 61.9 61.9
+ FasterVLM [56] 65.9 47.7 68.0 61.4 61.4
+ FreeVideoLLM [12] 55.5 36.0 68.9 59.3 57.6
+ iLLaVA [15] 64.5 56.6 65.5 61.5 59.3
+ LLaVA-PruMerge [42] 58.0 44.6 67.4 61.2 60.0
p-MoD 70.0 61.8 69.0 63.3 63.4
Table 10: Comparison with more vision token compression methods. In addition to the comparison in Table 5, we make a fair comparison between p-MoD and more vision token compression methods by controlling the average token retention ratio. Our methods achieves the best overall performance across 15 benchmarks. *Average is computed on all 15 benchmarks used in Table 1 and 3.
Model
Keep
[…]
Doc
VQA
@@ 15 Tasks*
Chart
QA
@@
RW
QA
SE
ED
PO
PE
$M$
MU
AI
2D
@@
$M$
E
G
QA
S
QA
$M$
B
AVG
LLaVA-v1.5 100% 28.1 25.8 18.2 46.0 55.6 66.2 85.9 36.6 55.2 76.6 53.4 1506.8 61.9 69.7 64.1 54.6
$\gamma-$MoD >  60% 20.4 21.5 18.1 47.5 53.7 66.3 86.6 35.0 55.1 77.1 51.3 1,377.0 62.1 67.2 59.9 52.7
+  p-MoD 53.7% 27.6 26.8 16.8 44.8 55.7 66.5 85.5 36.3 56.2 76.9 56.0 1482.8 62.2 69.3 65.4 54.7
Table 11: Comparison with concurrent work $\gamma$-MoD. All models are of 7B parameter scale.
@@
Appendix B Visualization of Token Selection Decisions
@@
Figure 6 visualizes the token selection decisions of p-MoD across different layers. The horizontal axis denotes the token indexes, and the vertical axis denotes the layer indexes. It can be observed that every layer select different tokens to process, and every token is selected by different p-MoD layers. This demonstrates that every p-MoD layer independently selects important and informative tokens, without degrading into selecting a same set of tokens across different layers, which is identical to dropping tokens instead of layer-wise selection.
Appendix C Further Discussions
C.1 Discussion on OCR Performance
In our experiments, we found that all token compression methods inevitably cause a significant performance drop on OCR-related benchmarks compared to other benchmarks. The main challenge we aim to address throughout the progress of this work is to mitigate this issue as much as possible. Experiment results in this paper show that our method achieves the least performance drop on OCR tasks compared to others.
C.2 Explanation on Training Overflow Problem
In Section 3.2.1, we mention that TanhNorm ensures training and inference stability. Acoordingly, we observe overflow issues in the ablation study on TanhNorm in Table 3. In this Section, we give a detailed explanation on the causes of overflow.
As illustrated in Equation 2, scaling a token by excessively large weights $w_{i}$ across multiple LLM layers may cause floating-point overflow. For TanhNorm with $\alpha=1$ (Table 3 Row 4), large $w_{i}$ makes $X_{i}^{\prime}=\alpha\tanh(w_{i})T(X_{i})+X_{i}$ approximate scaling the token $X_{i}$ by 2. Repeating this for the same token across multiple layers causes overflow. The same problem arises for vanilla MoD (Table 3 Row 1). In this case, no normalization is applied to constrain the range of the weights, making it easier to produce extreme values and result in numerical overflow.
C.3 Limitations
One limitation of our work is that p-MoD is only experimented on LLaVA-1.5 and LLaVA-NeXT models, which focus on single-image understanding tasks. We believe that our approach has the potential to achieve more remarkable results when applied on tasks that handle a larger number of vision tokens, such as multi-image and long video understanding. We leave the exploration of p-MoD on other vision tasks to future research.
Another limitation is that p-MoD is a trainable method, designed for training a new MLLM from scratch with reduced training and inference costs. When being applied to trained MLLMs, it requires continual fine-tuning the model. It is not suitable for training-free scenarios.
Appendix D More Implementation Details
D.1 Detailed Training Recipe